What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x + 5 through (2, 1)
Answer by ewatrrr(24785)

   
Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
line perpendicular to y = 1/3x + 5 through (2, 1)
Note: slopes of perpendicular lines are negative reciprocals of one another
New Line:
y = -3x + b ||using Pt(2,1) to solve for b
1 = -6 + b
7 = b
y = -3x + 7